OpenStudy (cleblanc17):

using these two equations y=10 and y=3/2x+1 how do i pair them into a system which can be solved using elimination or substitution?

OpenStudy (thomas5267):

\[ y=10\\ y=\frac{3}{2}x+1\\ 10=\frac{3}{2}x+1 \]

OpenStudy (cleblanc17):

lol y=10 is an equation .... @thomas5267

OpenStudy (thomas5267):

Yes! So?

OpenStudy (cleblanc17):

then how am i supposed to solve it by using substitution or elimination now? @thomas5267

OpenStudy (cleblanc17):

lol

OpenStudy (thomas5267):

You have substituted \(y=10\) into \(y=\dfrac{3}{2}x+1\) to get \(10=\dfrac{3}{2}x+1\). So just solve \(10=\dfrac{3}{2}x+1\) then put the x back into \(y=\dfrac{3}{2}x+1\).

OpenStudy (cleblanc17):

so i plug in 10=3/2x+1 into y???

OpenStudy (thomas5267):

Solve \(10=\dfrac{3}{2}x+1\) first. It is a linear equation with one variable.

OpenStudy (cleblanc17):